Q3. At 20 °C, the osmotic pressure of a polycarbonate was measured in chlorobenzene solution with the following results: Concentration (g/L) 1.95 2.93 3.915.86 Height of chlorobenzene (cm 0.20 0.36 0.53 0.98 Solvent density = 1.10 g/cm, polymer density = 1.20 g/cm3] Estimate the number average molecular weight Mn [Rz 8314 x 107 erg mol-1KY. (10 marks)
Answer #1

We will follow the following relation

π / c = RT / M + A2c

Where

π = osmotic pressure = height of chlorobenzene X density X acceleration due to gravity = hrg

g = acceleration due to gravity = 980 cm / second2

c = concentration (g / L)

R= gas constant = 8.314 X 107 erg / mol K

M = molar mass of polymer = ?

T = temperature = 293 K

A2= viral coeffecient

Therefore

\frac{h\rho g }{c}= \frac{RT}{M} + A_{2}c

Or

\frac{h }{c}= \frac{RT}{\rho gM} + \frac{A_{2}c}{\rho g}

so if we plot a graph between h/c v/s c we will get intercept = \frac{RT}{\rho gM}

From which we can calculate the molar mass

concentration (g/L) conce (g/mL) height h/c
1.95 0.00195 0.2 102.5641
2.93 0.00293 0.36 122.8669
3.91 0.00391 0.53 135.5499
5.86 0.00586 0.98 167.2355

h/c versus c 180 160 140 120 100 80 60 40 20 y= 16198x + 72.72 h/c versus c Linear (h/c versus c) 0.002 0.004 0.006 0.008

The equation is

y = 16198x + 72.72

Here intercept = 72.72 = \frac{RT}{\rho gM}

Molar mass = 8.314 X 107 X 293 / 1.10 X 9.8 X 102 X 72.72

Molar mass = 310745.61 g / mole